OpenStudy (anonymous):

1. Is (0,3) a solution to the equation y = x + 3? A. Yes B. No

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

put the value of y=3 and x=0 in the equation and then you find that it is correct

OpenStudy (texaschic101):

(0,3)...x = 0 and y = 3 now we sub y = x + 3 3 = 0 + 3 3 = 3 (correct) yes, it is

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

hurray you got the answer

OpenStudy (anonymous):

I'm sick so my brain doesn't work right haha

OpenStudy (anonymous):

but kelli keep in mind if this is on a question on a test dont cheat

OpenStudy (anonymous):

It isn't a test

OpenStudy (anonymous):

oh ok

OpenStudy (texaschic101):

hello kelli....my name is kellie also

OpenStudy (anonymous):

Really? Cool! :D

OpenStudy (anonymous):

wow wierd coincidence kelli it would be great if you would close the question

OpenStudy (texaschic101):

hope you get to feeling better :)

OpenStudy (anonymous):

Thank you :)